OpenStudy (ray1998):

Please help? Will give fan and medal. 1. Classify -6x^5 + 4x^3 + 3x^2 + 11 by degree. A. Quintic B. Cubic C. Quartic D. Quadratic 2. Write -2x^2(-5x^2 + 4x^3) in standard form. A. -8x^5 - 20x^4 B. -8x^5 + 10x^4 C. -7x + 2x^4 D. -7x^5 - 10x^4 Thank you in advance! :)

OpenStudy (ray1998):

BRB.

OpenStudy (anonymous):

http://www.cpsk12.org/cms/lib8/MO01909752/Centricity/Domain/3257/5.1%20ANSWERS.pdf may be this will help

OpenStudy (anonymous):

for Number 1, what is the largest degree?

OpenStudy (solomonzelman):

Degree of the polynomial is classified by the largest power. \(\sf Some~Examples: \) \(\color{#000000 }{ \displaystyle 3x^4+2x^3-7x+9 }\) The largest power/exponent in this equation is \(4\), therefore the polynomial is "quartic" polynomial. \(\color{#000000 }{ \displaystyle 3x^2-5x^5+2x }\) The largest power/exponent in this equation is \(5\), therefore the polynomial is "quantic" polynomial. \(\color{#000000 }{ \displaystyle x+4-7x^2+11x^3 }\) The largest power/exponent in this equation is \(3\), therefore the polynomial is "cubic" polynomial.

OpenStudy (solomonzelman):

What is the largest power/exponent in #1 ?

OpenStudy (ray1998):

Hey guys! Sorry, I'm back.

OpenStudy (ray1998):

exponent 5? @SolomonZelman

OpenStudy (anonymous):

I think @SolomonZelman meant "quintic" instead of "quantic", if that's what's confusing you.

OpenStudy (ray1998):

Okay, thank you :) @Alphabet_Sam

OpenStudy (anonymous):

every exponent has its own name

OpenStudy (anonymous):

On the second one, you just need to distribute the 2x^2 and then write it in order of descending exponent value. A more in-depth explanation of Standard Form for polynomials can be found here: http://www.softschools.com/math/algebra/topics/writing_polynomials_in_standard_form/

OpenStudy (ray1998):

Okay, thank you both so much! @SolomonZelman and @Alphabet_Sam

OpenStudy (ray1998):

What I should have done, was post the questions separately so that I can award you both. I'll repost so that both of you can receive a medal :) thank you for all your help.